Find the slope of AC and BD. Decide whether AC is perpendicular to BD.
A(-1,-2) B(-3,2) C(0,1) D(3,0)

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

Slope of Line Ac
x1 y1 x2 y2
-1 -2 0 1

slope m =(y2-y1)/(x2-x1)
(1-(-2)/(0-1)
(3/1)
m=3
Slope of line BD
x1 y1 x2 y2
-3 2 3 0

slope m =(y2-y1)/(x2-x1)
(0-2)/( 3-(-3) )
(-2/6)
m=-0.33
m1*m2=-1
so the lines are perpendicular


Answer by ikleyn(53862)

.
Find the slope of AC and BD. Decide whether AC is perpendicular to BD.
A(-1,-2) B(-3,2) C(0,1) D(3,0)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution in the post by @mananth is presented in monstrous form.
        He correctly determined that  m1 = 1/3,  but then he writes  m2 = -0.33.
        This is wrong.  What is  TRUE,  is that   m2 = -1/3,  but not  -0.33.
        So,  the correct presentation should be as showed below.


Slope of Line Ac
x1 y1 x2 y2
-1 -2 0 1
slope m1 =(y2-y1)/(x2-x1)
(1-(-2)/(0-1)
(3/1)
m1 = 3

Slope of line BD
x1 y1 x2 y2
-3 2 3 0
slope m2 =(y2-y1)/(x2-x1)
(0-2)/( 3-(-3) )
(-2/6)
m2 = -1/3
m1*m2=-1
so the lines are perpendicular


This is the correct form to present the solution.

----------------------

For @mananth,  -1/3  and  -0.33 is the same value,  but it is not true - they are  DIFFERENT !

In  Math,  to teach that  -0.33  is the same as  -1/3  - this is a  CRIME  ( ! )